### v1

@@ 4 Results
5. Automating Code Quality Checks: In SWSs, maintaining high code quality and consistency is essential for reliable data analysis and reproducibility. Tools like pre-commit ci are invaluable in this process, as they automate code quality checks and enforce coding standards. By running a series of predefined checks—such as linting, formatting, and security scans—automatically whenever code changes are made, pre-commit ci ensures that any issues are identified and addressed promptly. This mechanism, often called precommitci autoupdate, keeps the codebase aligned with the latest best practices and standards by automatically updating the checks. An example of this is seen in the pull request titled ’TS config and linting tweaks’, where pre-commit ci was used to enforce code quality standards. This pull request illustrates how automated checks can detect potential issues early in the development process, ensuring that the code adheres to the project’s guidelines before being merged. This pull request demonstrates how automated checks can catch issues early, ensuring the code adheres to the project’s guidelines before merging.
6. Managing Releases: Managing release is a crucial milestone that brings together various updates and improvements. When a new version is created, it signifies that a significant set of changes has been finalized and is ready for deployment. A generated release might follow, which involves automating the packaging and distribution of the updated software. Key elements of this process include addressing bug fixes to resolve issues reported by users and integrating features that enhance the software’s functionality. Additionally, a documentation release is often part of the process, ensuring users have up-to-date guides and references. Routine updates, or chore releases, might also be included to handle maintenance tasks such as dependency updates or configuration tweaks. For instance, ’Add note on fastapi to admin release notes’ exemplifies how significant updates and bug fixes are integrated into the software, reflecting the importance of a well-managed release process. Similarly, the issue ’Release Celery v5.3.5’ highlights how tracking and resolving bugs contribute to the overall quality and stability of the release.
7. User Interface: In SWSs, UI/UX design ensures that complex processes are accessible and manageable for researchers and scientists. A well-designed user interface (UI) and user experience (UX) can significantly enhance the usability of these systems, enabling users to seamlessly create, execute, and monitor workflows. Effective UI/UX in SWSs often involves intuitive layout and navigation, clear visualization of data and process steps, and interactive elements that facilitate user engagement and manipulation. By prioritizing user-centric design, these systems can reduce cognitive load, minimize errors, and improve the overall efficiency of scientific research. For example, features like drag-and-drop functionality for designing workflows or interactive dashboards for real-time monitoring can simplify intricate tasks and streamline workflows. Specific improvements such as ’Fix collection type drag and drop validation’ and ’Workflow invocation view improvements’ illustrate how targeted UI/UX enhancements can accelerate scientific discovery and innovation.
8. Improving Code Quality: In SWSs, maintaining high code quality is crucial for reliable and accurate computations. A common issue, known as the flake error, often arises from inconsistencies in code style or improper import order, as detected by tools like flake-import-order. For example, consider the issue ’Python package import collision’ in the Galaxy SWS. To resolve these problems, developers need to fix flake errors by adhering to consistent coding practices and arranging imports correctly. The xref mechanism is also vital, enabling the addition (xref add) and updating (xref update) of cross-references across modules and workflows, which helps track dependencies and maintain a clear structure. When an xref fix is necessary, correcting the error and updating the cross-references is important. Additionally, developers may need to remove outdated or conflicting components and pin specific library versions to ensure consistency and reproducibility. By implementing these practices, SWSs can become more reliable and maintainable, reducing errors and ensuring efficient execution of scientific computations.
9. Data Management and Task Scheduling: Effective data management and task scheduling are pivotal for optimizing performance and resource utilization in SWSs. The core concept revolves around breaking down complex tasks into smaller, manageable units, which are then assigned to workers by a scheduler. This scheduler uses sophisticated algorithms to determine the optimal sequence and allocation of tasks, ensuring efficient execution. One critical aspect is the partitioning of data, which involves dividing large datasets into smaller chunks or partitions. These partitions are distributed across workers, facilitating parallel processing and reducing the computational load on individual nodes. During the workflow, tasks may require data shuffling, where data is moved between workers based on specific criteria, ensuring that each worker has the necessary data to complete its assigned tasks. Data may be passed through the system’s memory or written to disk, depending on the workflow’s requirements and the system’s architecture. This topic discusses various aspects of effective scheduling and data management. For instance, one of the practitioners of SWSs raised an issue ’New jobs are not getting scheduled due to the error ”InternalError galaxy.jobs.handler in __handle_waiting_jobs’. Another pull request raised by other practitioners is ’[Dask.order] Ignore data tasks when ordering’.
10. Image Optimization: In SWSs, image optimization enhances performance and storage efficiency. Optimizing images involves reducing their size through compression techniques without significantly compromising quality. This process, known as image compression, can reduce the file size, making it easier to handle large datasets and speeding up data transfer and processing times. Image optimization techniques can also identify the more important image components and discard less important components. Additionally, improving filenames with descriptive and standardized formats ensures better organization and accessibility within the system. Tools like Calibre can be utilized for automated image actions, such as resizing and format conversion, further streamlining the workflow. By implementing these optimization strategies, scientific workflow systems can operate more efficiently, facilitating quicker data analysis and more effective resource management. One example of this topic is ’Auto Compress Images by Calibre’s image-actions’.
11. Annotation: In SWSs, annotation involves the addition of metadata or comments to various workflow components, such as data, processes, or results. These annotations can take the form of descriptions, tags, explanations, or contextual information, which collectively enhance the clarity, reproducibility, and usability of scientific workflows. This topic explores annotation issues, such as the problem where ’annotations somehow appear in workflow jobs’. In certain SWSs, like Celery, annotations play a significant role in adding extra information to tasks, thereby influencing their behavior or providing additional context. For example, the ‘max_retries‘ annotation in Celery specifies the maximum number of times a task should be retried if it fails. A specific issue related to this is ’Add annotations to minimize differences with celery-aio-pool’s tracer.py’
12. Kubernetes Deployments: In SWSs, efficient deployment and management of applications across Kubernetes clusters are essential. Helm, a powerful package manager for Kubernetes, plays a vital role in this process by using Helm charts to define, install, and upgrade complex applications. When setting up a new cluster, it’s essential to ensure that all nodes are configured correctly and Helm charts are deployed as intended. This topic discusses issues such as how to reproduce an issue reported by users and how to track and manage these effectively, from filing the initial report to potentially reopening it if the problem persists and eventually closing it once resolved. One such issue is ’Make helm chart compatible with $apache/nifi\geq 1.19.0$in Kubernetes cluster’.
13. Maintaining Changelog: Maintaining an accurate changelog is essential for any software project, including SWS. A detailed changelog provides a comprehensive record of all changes made in each release, helping both users and developers stay informed about new features, bug fixes, updates, or deprecations. It also includes documentation updates that clarify usage or functionality. Without a proper changelog, users might overlook important updates or changes, leading to confusion or issues during implementation. Therefore, consistently adding entries to the changelog for every release is a best practice, ensuring that all modifications are clearly documented. This topic addresses various issues related to changelogs, such as the problem of a missing changelog entry in an SWS project, exemplified by the issue ’Add missing changelog’.
Tables 2 and 3 present topics related to SWSs but from different sources—SOdata and GitHub. A notable similarity between the two tables is the presence of topics focused on data structures and operations, as well as task management and scheduling. The ’Data Structures and Operations’ topic in both tables highlights the significance of handling dataframes, arrays, and related data manipulation tasks within SWS. Additionally, task management is a recurring theme, with the ’Workflow Creation and Scheduling’ from SO data and ’Data Management and Task Scheduling’ from GitHub, both emphasizing the orchestration and execution of tasks within workflows. Moreover, both tables reference topics related to errors and bugs, with GitHub specifically addressing ’Errors and Bugs Fixing’, while SO topics such as ’Workflow Execution’ and ’Managing Rules, Input, Output, and Files’ implicitly involve error handling and troubleshooting. This overlap underscores the importance of managing data, tasks, and errors in SWSs across different platforms.
@@ 4.2 RQ2: What types of SWSs development questions are asked on technical forums?
Results: Table 4 shows the percentage of the posts types for each SWSs topic.
Figure 6: Types Distribution of SO Posts
The analysis of question distributions across various topics in SWSs reveals distinct patterns that underscore developers’ needs and knowledge gaps. The predominant focus on ’How’ questions, averaging 60.97% across all topics, indicates a significant demand for procedural guidance among users. For instance, topics such as ’Workflow Creation and Scheduling’ (51.75%) and ’Distributed Task Management’ (61.89%) demonstrate users’ strong interest in understanding the execution of specific tasks. Similarly, ’Processor for Data Processing’ (63.31%) and ’Data Structures and Operations’ (58.71%) highlight the necessity for detailed instructions and methodologies. Conversely, the ’Why’ questions, constituting an average of 17.2%, reflect users’ curiosity about the underlying reasons behind certain actions, with ’Workflow Execution’ and ’Task Dependencies and Management’ topics showing notable percentages (21.98% and 20.49%, respectively). ’What’ questions, accounting for 19.57% on average, indicate users’ need for clarification on the specifics of tasks or concepts, as seen in ’Workflow Execution’ (28.79%) and ’Automation with Ansible’ (24.68%). The ’Other’ category, with a minimal average of 2.26%, suggests that most user inquiries are well-categorized within the primary question types. Figure 6 shows the summary of types distribution for our analysis.
@@ 4.3 RQ3: Which topics are most difficult to answer?
Errors and Bugs Fixing 9.62 45.68
Documentation 7.18 36.33
Dependencies 0.15 51.60
Data Structures and Operations 19.78 105.21
Automating Code Quality Checks 0.76 0.20
Managing Releases 0.62 4.70
User Interface 7.86 17.19
Improving Code Quality 4.38 8.43
Data Management and Task Scheduling 12.45 61.44
Image Optimization 0.47 0.91
Annotations 13.53 22.86
Kubernetes Deployments 27.27 84.46
Maintaining Changelog 5.34 4.44
Table 7 shows the percentage of unresolved issues and pull requests and the median time required to resolve them. The analysis reveals that Data Structures and Operations and Kubernetes Deployments are among the most complex topics within scientific workflow systems on GitHub. With 19.78% of issues and pull requests unresolved and a median resolution time of 105.21 hours, Data Structures and Operations pose significant challenges, likely due to the intricate nature of managing large datasets and performing complex operations. Similarly, Kubernetes Deployments, with the highest unresolved rate of 27.27% and a median resolution time of 84.46 hours, highlight the complexity of deploying and managing applications in distributed environments, which require specialized expertise and thorough troubleshooting.
Another challenging area is Annotations, where 13.53% of issues remain unresolved, and the median time to resolve them is 22.86 hours. The difficulties here stem from the precision needed in implementing type validations and annotations, which are crucial for ensuring code quality but are often complicated and time-consuming to address. Data Management and Task Scheduling also ranks high in difficulty, with 12.45% unresolved issues and a median resolution time of 61.44 hours, reflecting the challenges in orchestrating and managing tasks within large-scale workflows.
On the other end of the spectrum, topics such as Automating Code Quality Checks and Image Optimization are relatively easy to manage, with very low percentages of unresolved issues (0.76% and 0.47%) and quick resolution times (0.20 and 0.91 hours). These tasks are well-supported by automation tools, making them straightforward to address. Managing Releases and Maintaining Changelog also fall into the easier category, benefiting from structured processes and automation that contribute to low unresolved rates and swift resolutions.
The analysis of difficulty across various SWSs topics on SO and GitHub reveals distinct patterns. On SO, topics like Workflow Execution (68.50% unresolved, 23.58 hours median time) and Workflow Creation and Scheduling (63.32%, 9.09 hours) have higher unresolved post percentages, indicating their complexity. In contrast, Data Transformation (34.08%, 1.81 hours) seems easier for developers. Similarly, on GitHub, Data Structures and Operations (19.78% unresolved, 105.21 hours) and Kubernetes Deployments (27.27%, 84.46 hours) present more challenges, with prolonged issue resolution times. Meanwhile, topics like Automating Code Quality Checks (0.76%, 0.20 hours) are resolved swiftly. This analysis highlights the diverse difficulty levels developers encounter in SWS, depending on the topic.
@@ 5.1 SWSs Topics Evolution
SWS is an emerging topic that is attracting increasing attention from developers across various domains, including software engineering [141], machine learning [142], bioinformatics [143], and data analysis[59]. To examine the evolution of SWSs, we present Figure 1, which shows the consistent growth of SWSs over time. The upward trend in the number of posts, issues, and pull requests indicates that SWSs topics are gaining significant attention from the community.
Figure 7: SWSs topics evolution over time on SO
To better understand the evolution of the different SWSs development activities, we measure the absolute growth of each of the ten topics over time shown in Figure 7. The data reveals a significant upward trend in the number of posts related to most topics, with ’Workflow Creation and Scheduling’ experiencing remarkably rapid growth, especially after 2015. This surge can be attributed to technological advancements that make workflow automation more accessible and essential for businesses. Tools and platforms like Apache Airflow, Kubernetes, and others have gained popularity, leading to more discussions. The proliferation of tools and platforms for workflow automation, coupled with a strong community focus on knowledge sharing, has further fueled this interest. Overall, the data underscores the critical importance of workflow and task management in enhancing productivity and efficiency in modern business environments.
@@
The absolute growth of GitHub 13 topics over time is shown in Figure 8. The evolution of topics related to SWSs reveals significant trends and shifts over the years. Initially, in 2009, the focus was primarily on documentation and error and bug fixing, with minimal attention to other areas. As the years progressed, there was a noticeable increase in the emphasis on ’Documentation’ and ’Errors and Bugs Fixing’, peaking around 2020. This indicates a growing recognition of the importance of thorough documentation and addressing errors in SWSs.
From 2015 onward, there was a substantial rise in topics related to ’Data Management and Task Scheduling’, ’Data Structures and Operations’, and ’Automating Code Quality Checks’, highlighting a growing focus on improving the efficiency and quality of SWSs. Notably, ’Data Structures and Operations’ saw a significant increase in attention, reflecting the importance of optimizing data handling within these systems. Similarly, the rise in ’Automating Code Quality Checks’ and ’Managing Releases’ demonstrates an evolving emphasis on maintaining and improving code quality and managing software releases. ’Image Optimization’ and ’Kubernetes Deployments’ emerged as relevant topics in later years, indicating the integration of new technologies and practices into SWSs. The data also show that ’User Interface’ and ’Improving Code Quality’ topics, while initially less prominent, have gained traction over time, underscoring the growing importance of user experience and code quality in SWSs.
Overall, the data reflects a broadening focus from basic documentation and error correction to more complex areas such as data management, code quality, and technological advancements, illustrating the evolving needs and priorities in the development of SWSs.

### v2

@@ 4 Results
12. Automating Code Quality Checks: In SWSs, maintaining high code quality and consistency is essential for reliable data analysis and reproducibility. Tools like pre-commit ci are invaluable in this process, as they automate code quality checks and enforce coding standards. By running a series of predefined checks, such as linting, formatting, and security scans, automatically whenever code changes are made, pre-commit ci ensures that any issues are identified and addressed promptly. This mechanism, often called precommitci autoupdate, keeps the codebase aligned with the latest best practices and standards by automatically updating the checks. An example of this is seen in the pull request titled ’TS config and linting tweaks’, where pre-commit ci was used to enforce code quality standards. This pull request illustrates how automated checks can detect potential issues early in the development process, ensuring that the code adheres to the project’s guidelines before being merged. This pull request demonstrates how automated checks can catch issues early, ensuring the code adheres to the project’s guidelines before merging.
This topic surfaces challenges with setting up and maintaining pre-commit hooks, CI pipelines, and style enforcement tools. Developers encounter integration friction between tools like flake8, pre-commit, and GitHub Actions, especially when working across teams with different environments. Misconfigurations and inconsistent enforcement reduce the effectiveness of automation.
13. System Redesign and API MigrationThis topic includes issues and pull requests related to large-scale system refactoring, API standardization, and migration efforts, often guided by formal proposals like AIP (Airflow Improvement Proposals). Keywords such as migrate, design, AIP47, example DAGs, and migration list highlight efforts to transition existing components (e.g., models, operators, and CI processes) to newer architectures or frameworks. For example, one issue notes: ’AIP44 Migrate Task Instance… and refactor local task join logic’, reflecting a migration that aims to decouple and improve task execution logic. These activities are essential for modernizing SWSs, ensuring scalability, maintainability, and alignment with long-term design goals.
The underlying cause of this challenge stemmed from the need to modernize legacy components while preserving compatibility and functionality. As SWSs evolve, older architectures and APIs often become inefficient, tightly coupled, or misaligned with new design principles (e.g., modularity, standardization). Migrating to updated frameworks, such as those introduced via formal improvement proposals (e.g., AIPs), requires careful refactoring, dependency tracking, and testing across complex components like DAGs, operators, and CI pipelines. These transitions are technically demanding and risky, often disrupted by incomplete migration paths, undocumented behaviors, or inconsistent integration patterns.
Tables 2 and 10 present topics related to SWSs but from different sources, Stack Overflow and GitHub. Based on the findings, the challenges developers face in SWS development are diverse and span both SO and GitHub platforms. The topic modeling analysis uncovered ten dominant topics on SO and thirteen on GitHub. On Stack Overflow, developers frequently struggle with issues like workflow creation and scheduling, distributed task management, and processor configuration for data processing, with challenges rooted in orchestrating tasks, managing DAGs, and debugging execution environments. GitHub, on the other hand, revealed additional pain points in areas such as error and bug fixing, documentation, and dependency management, which are critical during real-world SWS development and maintenance. Several topics, like data structures and operations, task coordination, and workflow scheduling, emerged as common across both platforms, suggesting a shared concern between learning and implementation contexts. These challenges reflect the complexity and multidisciplinary nature of SWS development, where developers must navigate domain-specific tools, distributed environments, and evolving community standards. Understanding these topics provides a foundational step toward designing tools and practices that better support SWS developers.
@@ 4.2 RQ2: What types of SWSs development questions are asked on technical forums?
Results: Table 14 shows the percentage of the posts types for each SWSs topic. The analysis of question distributions across various topics in SWSs reveals distinct patterns that underscore developers’ needs and knowledge gaps. The predominant focus on ’How’ questions, averaging 60.97% across all topics, indicates a significant demand for procedural guidance among users. For instance, topics such as ’Workflow Creation and Scheduling’ (51.75%) and ’Distributed Task Management’ (61.89%) demonstrate users’ strong interest in understanding the execution of specific tasks. Similarly, ’Processor for Data Processing’ (63.31%) and ’Data Structures and Operations’ (58.71%) highlight the necessity for detailed instructions and methodologies. Conversely, the ’Why’ questions, constituting an average of 17.2%, reflect users’ curiosity about the underlying reasons behind certain actions, with ’Workflow Execution’ and ’Task Dependencies and Management’ topics showing notable percentages (21.98% and 20.49%, respectively). ’What’ questions, accounting for 19.57% on average, indicate users’ need for clarification on the specifics of tasks or concepts, as seen in ’Workflow Execution’ (28.79%) and ’Automation with Ansible’ (24.68%). The ’Other’ category, with a minimal average of 2.26%, suggests that most user inquiries are well-categorized within the primary question types.
@@ 4.3 RQ3: Which topics are most difficult to answer?
Errors and Bugs Fixing 9.98 52.74
Documentation 8.03 40.50
Dependencies 0.02 44.62
Visual and Branding Enhancements 5.97 18.28
Automated Performance and Tool Update Integration 0.29 0.1
CRAN Release Management and Package Validation 4.37 85.53
Managing Releases 0.38 2.27
Browser Compatibility and HDFS Integration Issues 25.0 95.96
Chord Execution and Task Coordination Issues 9.7 91.45
Data Shuffling and Partitioning Strategies 10.12 75.17
Image Compression & Optimization 0.0 0.97
Automating Code Quality Checks 0.01 16.63
System Redesign and API Migration 3.85 117.95
Table 17 presents the proportion of unresolved issues and the median time to close issues across various identified topics in SWS development based on GitHub issue and pull request data. This analysis helps uncover persistent developer pain points and the relative difficulty associated with resolving them. Notably, Browser Compatibility and HDFS Integration Issues exhibit the highest rate of unresolved issues (25.0%) and a significantly long median resolution time (95.96 hours), indicating substantial complexity or lack of sufficient support in addressing such problems. Similarly, Chord Execution and Task Coordination Issues (9.7% unresolved, 91.45 hours) and Data Shuffling and Partitioning Strategies (10.12% unresolved, 75.17 hours) also show high-resolution times, suggesting that challenges involving distributed computation and task orchestration remain particularly difficult to resolve.
Conversely, topics such as Automated Performance and Tool Update Integration, Image Compression & Optimization, and Managing Releases show low unresolved rates and short median resolution times, indicating they are either well-supported or considered lower priority by the community. Interestingly, even topics like Documentation (8.03% unresolved) and Errors and Bugs Fixing (9.98% unresolved) have non-trivial unresolved rates, reinforcing that even foundational issues can persist over time. System Redesign and API Migration, while showing a moderate unresolved rate (3.85%), has the highest median resolution time (117.95 hours), possibly due to the complex, architectural nature of such tasks. Overall, the data highlights that unresolved rates and resolution times vary significantly by topic, with issues related to distributed execution, integration, and large-scale redesign emerging as particularly burdensome for developers. These findings point to critical areas where improved tooling, documentation, or community support could help alleviate recurring developer challenges in SWS development.
The analysis of difficulty across various SWSs topics on SO and GitHub reveals distinct patterns. On SO, topics like Workflow Execution (68.50% unresolved, 23.58 hours median time) and Workflow Creation and Scheduling (63.32%, 9.09 hours) have higher unresolved post percentages, indicating their complexity. In contrast, Data Transformation (34.08%, 1.81 hours) seems easier for developers. Similarly, the GitHub analysis revealed that System Redesign and API Migration had the highest median time to close issues, at 117.95 hours, followed by Browser Compatibility and HDFS Integration Issues (95.96 hours) and Chord Execution and Task Coordination (91.45 hours). Notably, Browser Compatibility and HDFS Integration also exhibited the highest proportion of unresolved issues (25.0%), highlighting substantial challenges related to system maintenance and cross-platform compatibility. Other topics, such as Data Shuffling and Partitioning Strategies and CRAN Release Management, also demonstrated prolonged resolution times, reflecting persistent developer struggles with performance tuning, dependency handling, and validation workflows within SWS environments. In contrast, topics like Image Compression and Optimization and Automated Performance and Tool Update Integration showed median resolution times below one hour and near-zero unresolved rates, suggesting these areas pose minimal difficulty or are well-supported by existing tooling. Overall, these findings indicate that developers face the greatest difficulty in infrastructure-level orchestration, fault-tolerant execution, and managing breaking system-level changes. Addressing these pain points through enhanced automation, robust deployment support, and improved documentation is essential to reduce the development burden and improve the overall maintainability of SWSs
@@ 5 Discussion & Implications of our Findings
Figure 6: SWSs topics evolution over time on SO
@@ 5.1 SWSs Topics Evolution
SWS is an emerging topic that is attracting increasing attention from developers across various domains, including software engineering [147], machine learning [148], bioinformatics [149], and data analysis[61]. To examine the evolution of SWSs, we presented Figure 1, which shows the consistent growth of SWSs over time. The upward trend in the number of posts, issues, and pull requests indicates that SWSs topics are gaining significant attention from the community.
To analyze how developer challenges in SWSs have evolved over time, we examined the year-wise distribution of all identified topics across both SO (Fig: 6) and GitHub (Fig: 7). On SO, early years were dominated by foundational concerns such as Workflow Creation and Scheduling and Distributed Task Management, which continued to grow steadily over the past decade. These trends reflect ongoing difficulties in defining pipelines, managing execution orders, and integrating workflow engines like Airflow and Nextflow. In recent years, the rise of topics such as Processor for Data Processing, Workflow Execution, and Task Dependencies and Management indicates a shift toward runtime optimization, resource allocation, and fine-grained control over pipeline execution.
@@
In parallel, GitHub exhibited a different yet complementary evolution. Errors and Bugs Fixing remained the most persistent topic across all years, underscoring its centrality in practical software maintenance. Documentation also showed a notable increase, reflecting the community’s growing attention to tool usability, onboarding, and reproducibility. Emerging topics like Workflow Execution and Processor for Data Processing mirrored the recent SO trends, while GitHub uniquely highlighted platform-specific challenges such as Browser Compatibility and HDFS Integration Issues and Chord Execution and Task Coordination. These reflect low-level system dependencies and orchestration concerns that are more likely to surface in code-level issue tracking than in general developer Q&A forums. Overall, Stack Overflow topics tend to center on conceptual understanding, best practices, and troubleshooting common errors, whereas GitHub discussions more frequently capture in-depth implementation, system-specific debugging, and long-term maintenance issues. The alignment and divergence of topic trajectories across platforms reveal the layered nature of SWS development, from initial learning and design on SO to deployment and sustainability on GitHub.
Figure 8: Overall evolution of SWSs Data
Figure 8 presents the evolution of overall topic-related developer activity on SO and GitHub over time. The number of GitHub issues and pull requests shows a marked increase beginning in 2015, reflecting the growing use of GitHub for collaborative development and issue tracking in SWSs. In contrast, SO activity gradually rose until it peaked around 2020, after which it plateaued or slightly declined. This divergence may indicate that while SO remains a valuable platform for asking general and technical questions, GitHub has increasingly become the central hub for hands-on debugging, issue resolution, and workflow collaboration. The visualization underscores the temporal dynamics of developer engagement and complements our topical analysis by showing how platform usage and topic discussion volume have shifted over time.